OpenStudy (anonymous):

How do you multiply this? (y+5) (y-1)

OpenStudy (lgbasallote):

do you know distributive property?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

then change this into \[y(y+1) + 5(y+1)\] use distributive on both terms

OpenStudy (lgbasallote):

that should be y-1 but you get the point ;)

OpenStudy (anonymous):

(y+5)(y-1)= y^2-1y+5y-5= =y^2+4y-5

OpenStudy (anonymous):

thank you

OpenStudy (asnaseer):

you can use the FOIL method. this site might help you understand this better: http://www.algebrahelp.com/lessons/simplifying/foilmethod/